For the following equation, state the value of the discriminant and then describe the nature of the solutions:
18x^2+5x-9=0
1. What is the value of the discriminant?
2. Are there two imaginary solutions?
3. Are there two real solutions?
4. Only one real solution?
Thanks for the help!

Answer by stanbon(75887)

For the following equation, state the value of the discriminant and then describe the nature of the solutions:
18x^2+5x-9=0
1. What is the value of the discriminant?
b^2-4ac = 5^2 - 4*18*-9 > 0
=====================================
2. Are there two imaginary solutions? no
----------------------------------------
3. Are there two real solutions? yes
----------------------------------------
4. Only one real solution? no
==================================
Cheers,
Stan H.